OpenStudy (anonymous):

\[ \frac{x^{4}}{x^{2}}. \] That was part 1. Part 2 says, Create your own fraction with like bases, coefficients, and show its simplification. Can someone help me please.

OpenStudy (lgbasallote):

hmm...name a letter...any of the english alphabet...

OpenStudy (anonymous):

x

OpenStudy (lgbasallote):

now give me a number

OpenStudy (anonymous):

1

OpenStudy (lgbasallote):

not 1...or zero

OpenStudy (anonymous):

7

OpenStudy (anonymous):

7x

OpenStudy (lgbasallote):

good good...so we'll use x as base and 7 as coeficient...it says fraction wherein the numerator and denominator has these so it;s \[\frac{7x}{7x}\] BUT WAIT! we add an exponent to x...now give me two different numbers

OpenStudy (anonymous):

2 and 3

OpenStudy (lgbasallote):

ok good 2 and 3...so it's \[\large \frac{7x^2}{7x^3}\] got it?

OpenStudy (anonymous):

Yes.

OpenStudy (lgbasallote):

to simplify that...we just cancel the coefficients since they are the same \[\frac{\cancel{7}x^2}{\cancel{7}x^3}\] got it?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

That part gets me.

OpenStudy (anonymous):

Why do you have to cancel them out?

OpenStudy (lgbasallote):

look at it this way \[\large \frac{7x^2}{7x^3} = \frac{7}{7} \times \frac{x^2}{x^3}\] okay?

OpenStudy (anonymous):

How would I even read that? Well, I kinda get it. but why cross out?

OpenStudy (anonymous):

Brb one sec. Need water.

OpenStudy (lgbasallote):

that's 7/7 times x^2/x^3 lol and it would cross out because 7/7 is 1 got it?

OpenStudy (anonymous):

Ohhhhhhhhhhh.

OpenStudy (anonymous):

I get it now.

OpenStudy (anonymous):

So, there is no more 7??

OpenStudy (anonymous):

At all in the expression/equation?

OpenStudy (lgbasallote):

yup just good old \(\Large \frac{x^2}{x^3}\)

OpenStudy (anonymous):

.. so why even have numbers? ....

OpenStudy (lgbasallote):

do you know how to simplify that?

OpenStudy (anonymous):

I do, but I don't. :/

OpenStudy (anonymous):

x^-1

OpenStudy (anonymous):

I'm unsure. :/

OpenStudy (lgbasallote):

yup that's right! x^-1..but how would you make it positive exponent?

OpenStudy (anonymous):

Umm. Take the negative out and add a one as the numberator and have x^1 as denom.

OpenStudy (anonymous):

numberator? lol. numerator.

OpenStudy (lgbasallote):

right! the simplified form is \[\frac{1}{x}\]

OpenStudy (anonymous):

x

OpenStudy (anonymous):

1 over x or just x? is there a differenec?

OpenStudy (lgbasallote):

let's say x is 2...is 1 over 2 the same as just 2?

OpenStudy (anonymous):

No.

OpenStudy (lgbasallote):

im not saying x is 2...im just showing it..

OpenStudy (lgbasallote):

so then 1/x is not x :DD

OpenStudy (anonymous):

lol. :) So to bring this to a summarization type deal.. how would I explain what we just did, a lot easier and/or in one message?

OpenStudy (anonymous):

Can we do another example? or another problem or whatnot?

OpenStudy (lgbasallote):

sure...try giving an example and ill help you simplify it

OpenStudy (anonymous):

Okay. Let me smoke this cigarette and I'll come back. Will you still be here to help me?

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

Alright. Give me a few minutes okay? Thank y9ou by the way.

OpenStudy (anonymous):

I'm back.